# On the Intrinsic Self-Correction Capability of LLMs: Uncertainty and Latent Concept

### v1

@@ Abstract
Large Language Models (LLMs) can improve their responses when instructed to do so, a capability known as self-correction. When these instructions lack specific details about the issues in the response, this is referred to as leveraging the intrinsic self-correction capability. The empirical success of self-correction can be found in various applications, e.g., text detoxification and social bias mitigation. However, leveraging this self-correction capability may not always be effective, as it has the potential to revise an initially correct response into an incorrect one. In this paper, we endeavor to understand how and why leveraging the self-correction capability is effective. We identify that appropriate instructions can guide LLMs to a convergence state, wherein additional self-correction steps do not yield further performance improvements. We empirically demonstrate that model uncertainty and activated latent concepts jointly characterize the effectiveness of self-correction. Furthermore, we provide a mathematical formulation indicating that the activated latent concept drives the convergence of the model uncertainty and self-correction performance. Our analysis can also be generalized to the self-correction behaviors observed in Vision-Language Models (VLMs). Moreover, we highlight that task-agnostic debiasing can benefit from our principle in terms of selecting effective fine-tuning samples. Such initial success demonstrates the potential extensibility for better instruction tuning and safety alignment. Warning: examples in this paper contain offensive language.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) research by contributing to state-of-the-art results for various downstream applications [1, 2, 3]. Despite the significant achievement of LLMs, they can generate harmful contents, e.g., toxicity [4, 5], harmfulness [6, 7], and bias [8, 9]. The primary reason is that LLMs are pre-trained on corpora collected from the Internet, wherein stereotypical, toxic, and harmful content is relatively common. In the pre-training stage, the safety alignment technique [10, 11] has become the de-facto solution for mitigating safety issues. However, safety alignment is not perfectly robust [12, 13, 14], remaining failure cases [6, 8] in practice. The recently proposed self-refine pipeline [15] stands out as an effective solution, leveraging the intrinsic self-correction capability of LLMs to improve performance with instruction prompting rather than computationally expensive fine-tuning. The self-correction pipeline [^1] In this paper, self-correction refers to both the intrinsic self-correction capability and the pipeline for leveraging the self-correction capability. only requires specific instruction designs to guide the LLM toward desired responses, self-correction instructions can be either directly concatenated with the original prompt or be appended to LLMs’ responses as the post-hoc prompt, to fix problems in previous responses. Self-correction has been widely adopted in many other applications including improving translation quality [16], defense jailbreak attack [17], optimizing code readability [15], among others.
Though there are empirical successes and studies of self-correction, its effectiveness remains a mystery and is under scrutiny [18, 19, 20]. One failure case is improper self-correction instruction design, [20] observes the failure of the critical self-correction instruction. In contrast, [21] proposes IoE instruction considering the confidence characteristic of LLMs, achieving consistent empirical success. In this paper, we elucidate the effectiveness of the intrinsic self-correction capability by characterizing how model uncertainty and the latent concepts activated by self-correction instructions contribute to the convergence of self-correction processes. Previous research has utilized calibrated uncertainty [22, 23, 24] to determine whether LLMs possess the internal knowledge necessary to correctly answer a given question [22]. Uncertainty has also been employed in active learning [25] and core-set selection [26]. Latent concepts have proven to be a practical feature aiding in understanding the mechanisms of in-context learning [27, 28]. The latent concept in this paper is defined as underlying intent of a text, e.g., bias or toxicity [29]. One example is the surgeon asked the nurse a question, he, wherein the statement deliveries an implicit gender stereotype that surgeons should be male.
Present work: we investigate how and why self-correction can be effective by the lenses of model uncertainty and activated latent concepts. Our study proceeds as follows: First, we show empirical evidence that, given properly self-correction instructions and a multi-round question-answering scenario, LLMs can perform better as the round increases and can reach a converged performance after several rounds. The convergence phenomenon happens to all considered tasks. Second, we investigate the underlying reason why the LLM can reach the convergence status. Both empirical evidence and theoretical analysis advocate the collaborative effect of model uncertainty and the activated latent concept by the instruction. The latent concept is the force driving uncertainty and self-correction towards convergence. Notably, our conclusions can also be extended to the self-correction behaviors observed in Vision-Language Models (VLM). Third, We demonstrate the feasibility of these insights to provide guidance for practical applications. We showcase an application to appropriately select task-agnostic debiasing samples for safe and robust social bias mitigation, with the goal of preserving the language modeling ability to the greatest extent possible. Empirical evidence indicates the selected samples can reduce the diminishment of language modeling ability. Such initial success also demonstrates the potential to extend the sample-selection criteria for selecting good instructions or fine-tuning samples, e.g., self-instruct [30] and self-alignment [31].
2 Main Analysis
@@
In this section, we explore the effectiveness of self-correction with comprehensive analyses of different tasks. The adopted tasks can be roughly categorized into (1) multiple-choice tasks including social bias mitigation [8], jailbreak defense [17], and visual question answer (VQA) (2) generation tasks including commonsense generation [32], text detoxification [4, 33], and visual grounding. Notably, visual grounding and visual question answer (VQA) are multi-modality tasks requiring an understanding of both vision and language. The model we majorly utilized is zephyr-7b-sft-full [34], an LLM model further fine-tuned on Mistral-7B-v0.1 [35] with instruction-tuning. GPT-4  [^2] https://openai.com/index/gpt-4-research/ is utilized as the backbone vision-language model for vision-language tasks. The self-correction instructions are utilized per round. The instruction for the first round is concatenated with the original question. The following instructions are appended with the dialogue history as the post-hoc instruction to correct the misbehavior. Following the setting in [20], we set the number of post-hoc instructions as a constant rather than using the correct label to determine when to stop. More experimental details can be found in Appendix D.
Experimental results of self-correction on different tasks can be found in Figure 1 where the x-axis represents the round of the instructions and the y-axis indicates the performance evaluated on the corresponding task. More experimental setting details can be found in Appendix D, respectively. Accordingly, we can have the following observations: Obs.1: Self-correction generally leads to a better performance compared to the baseline instruction wherein no self-correction purpose is incorporated. Obs.2: Multi-round self-correction can guide LLMs to a convergence state with a stable performance. The response does not further change with additional self-correction steps. Obs.3: Multiple-answer choice problems converge in the first iteration while generation tasks require more rounds toward the final convergence. A potential reason is that free-form text generation is generally more complicated than a closed-form choice selection task. In conclusion, the self-correction with appropriate instruction leads to consistent performance gains and convergence. Those observations motivate an investigation into the underlying mechanism that enables LLMs to achieve a convergence state with satisfying performance.
@@
Figure 1: The self-correction performance for six different tasks including both generation and multiple-choice tasks. The x-axis represents the iteration round and the y-axis indicates the performance evaluated on the corresponding task. The performance of self-correction improves as the interaction round progresses and converges eventually. The self-correction performance of the social bias mitigation task and the jailbreak defense task reaches the best in the first round and keeps the best performance with no modification for the rest interaction rounds.
Organization. In light of the above observations, we endeavor to understand the underlying reasons for the above observations in the following sections. Section 2.1 elucidates that the decreasing model uncertainty is one factor correlated with the convergence state of self-correction. Section 2.2 unveils another key reason for the convergence is the self-correction instruction can help to activate the task-aware positive concept, e.g., harmless, fairness. Section 2.3 shows how model uncertainty and the activated positive concepts correlated to the performance of self-correction. In addition to comprehensive empirical analyses, Section 2.4 presents a more elegant mathematical formulation for self-correction, underscoring the driving force of activated concept to the convergence of both uncertainty and self-correction. Theoretical analyses also shed light on the significant role of the first-round instruction for the final converged performance, this is aligned with our empirical observations.
Our analyses can provide practical suggestions for self-correction as follows: (1) Iteratively applying a proper self-correction instruction can lead to a satisfied and converged performance. (2) An appropriate first-round instruction is of most effective to performance improvement with the least computational cost. (3) Multiple-choice tasks only require one round of instruction with no further improvement from more self-correction steps, though the activated concept and mode uncertainty evolve incrementally better. (4) Better self-correction instructions enjoy lower uncertainty.
2.1 Decreasing Model Uncertainty
Based on previous observations where the self-correction performance converges without further improvement, we hypothesize that one underlying reason is the decreasing uncertainty upon interaction round progresses. Uncertainty typically reflects the model’s internal assessment on the reliability of its own responses. To estimate the uncertainty, we adopt the method of semantic entropy [36], which involves estimating linguistic-invariant likelihoods by the lens of semantic meanings of the text rather than regarding input texts equally. We estimate the uncertainty of LLM’s response on each round with experimental results shown in Figure 2.
(a) Generation tasks
(b) Multiple-answer choice tasks
Figure 2: The reported uncertainty of four tasks across interaction rounds. For language generation tasks (a), we interact with the LLM for 10 rounds, and the interaction rounds for multi-answer choice tasks (b) is 5. All the reported tasks show a tendency of convergence in terms of uncertainty as the round progresses. The uncertainty among generation tasks (a) declines to a stable status as the interaction round increases, but the uncertainty for multi-answer choice tasks (b) generally approaches the lowest level in the first round, consistent with its performance with no modification.
We can then have the following observations: Obs.1: The uncertainty generally decreases along with more self-correction rounds across tasks. Obs.2: All the reported tasks demonstrate a trend of converging uncertainty as the rounds progress. Obs.3: Multiple-choice tasks demonstrate low uncertainty from the first iteration, which helps to explain why the LLM converges in the first iteration. Obs.4: The LLM can be more certain on the multiple-choice tasks as it only requires to compare within given restricted choices. To sum up, self-correction instructions can persistently decrease the answer uncertainty, leading to an invariant final response. Nonetheless, it remains unclear why the response generated with self-correction instructions exhibits better performance than that generated with a baseline instruction.
2.2 Activated Latent Concept
In this section, we investigate how the activated latent concept evolves as the self-correction process progresses and drives performance improvement, building on the approach of identifying latent concepts to understand in-context learning [27, 28]. We hypothesize that self-correction instructions can activate task-aware latent concepts (henceforth referred to as activated concepts), thereby enhancing the performance of large language models. The term latent concept has been employed to explore underlying mechanisms of in-context learning [27, 28, 29]. However, there is no consensus on the precise definition of the concept. For instance, [27] considers the transition probability of an HMM-based language generation model as the concept, while  [28] defines it as the latent topic of a given text. In this context, a latent concept is regarded as a task-aware latent intent underlying the input. For example, in the social bias mitigation task, the negative concept corresponds to bias, whereas the positive concept represents fairness. Similarly, in the text detoxification task, the concepts include toxicity and non-toxicity.
To empirically validate our hypothesis, we utilize typical probing vectors for each task and measure the similarity between these probing vectors and the hidden state representations of input sequences. The probing vector [37] only incorporates the information relevant to one latent concept e.g., toxicity. Measuring the similarity between a hidden state and a probing vector can guarantee the hidden state is projected to a clean space associated with the target latent concept. We employ a typical strategy to acquire the probing vector for each task: (1) social bias mitigation: rephrasing input samples to exhibit bias or harmfulness; (2) jailbreak defense: discovering the latent knowledge of concepts through CCS [38]; (3) text detoxification: probing the hidden state representation with a trained classification model. For the VAQ and visual grounding task, we use the GPT-4 as the backbone model so we can not access the hidden state representation. With respect to the commonsense generation task, the latent concept should be very complex and contain various linguistic features. Therefore we only analyze the latent concept for the tasks of social bias mitigation, text detoxicification, and jailbreak defense. More details about probing vector acquisition are available in Appendix D.4.
Figure 3: The similarity between the activated latent concept and the associated probing vector of task-aware (positive) concepts was examined across three tasks. Higher similarity values indicate that a more task-aware latent concept has been activated. All tasks exhibit a trend of increasing similarity followed by a slight decline.
The similarity between the activated latent concept and the probing vector across interaction rounds is presented in Figure 3 . The overall trend observed across these tasks initially shows an increase, followed by a subsequent decline. It is evident that task-aware concepts are activated during interactions, with the second round contributing to the most significant increase. For the subsequent decline phenomenon, we posit that the underlying reason is that the dialogue history may hinder the activation of positive concepts. This phenomenon is further analyzed in our theoretical discussion in Section 2.4. In summary, self-correction instructions facilitate the activation of positive concepts across interaction rounds. However, the activated positive concept does not remain consistent and exhibits a slight decline over time.
2.3 The Collaborative Effect of Uncertainty and Latent Concept

### v2

@@ Abstract
Warning: examples in this paper contain offensive language.
Large Language Models (LLMs) are able to improve their responses when instructed to do so, a capability known as self-correction. When instructions provide only the task’s goal without specific details about potential issues in the response, LLMs must rely on their internal knowledge to improve response quality, a process referred to as intrinsic self-correction. The empirical success of intrinsic self-correction is evident in various applications, but how and why it is effective remains unknown. In this paper, we unveil that intrinsic self-correction can be progressively improved, allowing it to approach a converged state. Our findings are verified in: (1) the scenario of multi-round question answering, by comprehensively demonstrating that intrinsic self-correction can progressively introduce performance gains through iterative interactions, ultimately converging to stable performance; and (2) the context of intrinsic self-correction for enhanced morality, in which we provide empirical evidence that iteratively applying instructions reduces model uncertainty towards convergence, which then leads to convergence of both the calibration error and self-correction performance, ultimately resulting in a stable state of intrinsic self-correction. Furthermore, we introduce a mathematical formulation and a simulation task indicating that the latent concepts activated by self-correction instructions drive the reduction of model uncertainty. Based on our experimental results and analysis of the convergence of intrinsic self-correction, we reveal its underlying mechanism: consistent injected instructions reduce model uncertainty which yields converged, improved performance.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing research by contributing to state-of-the-art results for various downstream applications [1, 2, 3]. Despite the significant achievements of LLMs, they are known to generate harmful content [4, 5], e.g., toxicity [6, 7] and bias [8, 9] in text. The primary reason for this is that LLMs are pre-trained on corpora collected from the Internet, wherein stereotypical, toxic, and harmful content is common. Thus, safety alignment techniques [10, 11] have become the de-facto solution for mitigating safety issues. However, safety alignment is not perfectly robust [12, 13, 14, 4, 8].
The recently proposed self-refine pipeline of  madaan2023self [15] stands out as an effective solution, leveraging the self-correction capability of LLMs to improve performance by injecting self-correction instructions or external feedback into the prompt. The self-correction pipeline [^1] In this paper, self-correction refers to both the self-correction capability and the pipeline for leveraging the self-correction capability. only requires specific instructions designed to guide the LLM towards desired responses; to correct errors in previous responses, these self-correction instructions can be either directly concatenated with the original prompt or appended to the LLMs’ responses as a post-hoc prompt. Self-correction has been widely adopted in many other applications, including improving translation quality [16], defense against jailbreak attacks [17], and optimizing code readability [15].
Intrinsic self-correction, as highlighted by [18], emerges as a more efficient method, as it does not require costly feedback from humans or more advanced LLMs. Instead, it relies solely on the model’s internal knowledge to address issues in responses. Furthermore, the instruction for intrinsic self-correction is very abstract and simple, such as Please do not be biased or rely on stereotypes. This example instruction directly describes the task-wise objective for the purpose of self-correction and does not deliver any specific details about the LLMs’ responses.
Though the empirical success of intrinsic self-correction across various applications has been shown [^2] Notably, in this paper, we omit consideration of reasoning tasks due to the existing debate about the effectiveness of self-correction for reasoning [19]., its effectiveness remains a mystery [20, 21, 19, 22]. There are two main research questions concerning intrinsic self-correction: (1) Can we guarantee that we can achieve convergence by iteratively applying intrinsic self-correction? This convergence guarantee is a fundamental prerequisite for practical utilization of the intrinsic self-correction capability. (2) What is the underlying reason for this convergence, if it exists? To answer these research questions, we explore the task of moral self-correction, analyzed through Question Answering (QA) scenarios [18], since the relevance of morality is one of the most important challenges to overcome when leveraging LLMs.
Figure 1 illustrates how we leverage the common setting of self-correction in a multi-round QA scenario to investigate how the latent concept and model uncertainty help the convergence in order to improve text detoxification performance. Model uncertainty has been utilized to quantify confidence levels in LLM predictions. [23, 24, 25, 26]. In this paper, we define the latent concept as the underlying moral orientation of an input text, e.g., latent stereotypes or toxic language underlying or implied by the text. One example is the surgeon asked the nurse a question, he, wherein the statement expresses an implicit gender stereotype that surgeons should be male. Latent concepts that have been activated by instructions have been proven to be a critical signal in the mechanistic understanding of in-context learning [27, 28] and morality in LLMs [29, 12]. In this work, we show that through multi-round intrinsic self-correction, latent concepts are activated, which reduces model uncertainty towards convergence, thus improving overall LLM performance.
Figure 1: Applying multi-round intrinsic self-correction for the task of text detoxification in a question-answering scenario. By injecting self-correction instructions (bold font) into queries (green text boxes) for several rounds, the toxicity level of generated sentences (blue text boxes) decline and ultimately approach convergence. Our experiments show this convergence can be achieved, on average, within 6 rounds of self-correction. We investigate how the latent concept and model uncertainty drive LLMs towards  convergence, thus achieving stable performance on downstream tasks, e.g., decreasing toxicity. By injecting instructions during multi-round self-correction, concepts are activated and model uncertainty is reduced.
Summary. By investigating LLMs’ intrinsic self-correction behaviors on morality-related tasks, our analysis shows that rounds of self-correction instructions reduce model uncertainty, which leads to convergence in calibration errors, ultimately resulting in stable performance of intrinsic self-correction on downstream tasks. The convergence and effectiveness of intrinsic self-correction directly arises from this reduction in uncertainty.
Organization. Section 2 presents the motivation for our hypothesis that intrinsic self-correction instructions reduce calibration errors by decreasing model uncertainty, driving the model towards converged performance. Section 3 shows empirical evidence that the convergence guarantee exists for various tasks. Section 4 elucidates how intrinsic self-correction reduces model uncertainty, i.e., a reduction in calibration error, until convergence of the calibration error. Section 5 illustrates how the activated latent concept evolves through self-correction rounds. Section 6 highlights the role of activated latent concepts as a driving force behind the convergence of self-correction performance, both empirically and theoretically.
2 Preliminary & Motivations
Background. In the context of machine learning, model uncertainty reflects how confident a model is in its predictions or generations [30, 31, 25]. For classification tasks, uncertainty is often quantified through prediction logit confidence [32]. However, in language generation tasks, the definition of uncertainty remains a topic of debate, with proposals ranging from verbal confidence [33] to semantic uncertainty [34]. In this paper, we adopt semantic uncertainty as the model uncertainty estimator for language generation tasks. For QA tasks, we reformulate them as classification problems by normalizing logits over the negative log-likelihood of each choice.
Previous studies demonstrate that avoiding over-confident or under-confident predictions can achieve calibrated uncertainty [35, 36]. Calibrated uncertainty characterizes to what extent LLMs’ prediction confidence aligns to the actual accuracy of those predictions [37, 24]. In our experiments, we show that LLMs are initially under-confident (high uncertainty) without the self-correction instructions. If a model is well-calibrated, its prediction confidence reflects the actual accuracy of those predictions. Therefore, the level of calibration error can be used to determine whether we can trust a prediction. In the context of LLMs, smaller calibration errors indicate that LLMs are more confident that they can answer the given question correctly, thereby, it also demonstrates better performance [23].
Figure 2 shows the logical framework of our analysis to reveal the convergence nature of intrinsic self-correction. We hypothesize that intrinsic self-correction effectively reduces model uncertainty by enhancing prediction confidence in QA tasks and minimizing semantic variability in language generation tasks. This reduction in uncertainty is achieved by incorporating self-correction instructions, which activate appropriate latent concepts [27]. Here, we define latent concepts as the underlying moral orientation within an input sentence [12], such as toxicity or implied stereotypes. Additionally, we provide both empirical and mathematical evidence demonstrating the dependence between model uncertainty and latent concepts. This establishes a logical progression from self-correction instructions (via latent concepts) to reduced model uncertainty, leading to lower calibration error and ultimately improved self-correction performance.
Figure 2: The logical framework of our analysis considers two key variables: the concept and model uncertainty. A positive concept implies that the activated concept aligns with the self-correction objective, such as fairness or non-toxicity. We hypothesize that the injected self-correction instruction can activate the desired concept, which in turn reduces model uncertainty. This reduction in model uncertainty is expected to decrease and stabilize the calibration error, ultimately leading to converged self-correction performance.
[…]
$p(y_{t}|q_{t},\theta)=\sum_{c\in\left[C_{p},C_{n}\right]}p(y_{t}|c,q_{t},\theta)\underbrace{p(c|q_{t},\theta)}_{\text{{latent concept}}}$ (1)
Equation 1 theoretically demonstrates the relationship between the latent concept, activated by the input $q_{t}$, and model uncertainty. To ensure that $q_{t}$ keeps activating $C_{p}$ across rounds, in Section 5 we empirically demonstrate that, by injecting proper instructions, the activated concept is not revertable.
3 The General Convergence of Intrinsic Self-Correction
@@
Experimental Settings. The adopted tasks can be categorized into (1) multi-choice QA tasks: social bias mitigation [8], jailbreak defense [17], and visual question answer (VQA) (2) generation tasks: commonsense generation [38], text detoxification [6, 39], and visual grounding [40]. Notably, visual grounding and visual question answer (VQA) are multi-modality tasks requiring an understanding of both vision and language. The considered model in this paper is zephyr-7b-sft-full [41], a LLM model further fine-tuned on Mistral-7B-v0.1 [42] with instruction-tuning. GPT-4  [^4] https://openai.com/index/gpt-4-research/ is utilized as the backbone vision-language model for vision-language tasks. We consider a multi-round self-correction pipeline in a QA scenario, and self-correction instructions are utilized per round. The instruction for the first round is concatenated with the original question. The following instructions are appended with the dialogue history as the post-hoc instruction to correct the misbehavior. Following the setting in huang2023large [19], we set the number of self-correction rounds as a constant rather than using the correct label to determine when to stop. We use 10 rounds for text detoxification and commonsense generation, and 5 rounds for other tasks. More experimental details can be found in Appendix B.
@@
Figure 3: The self-correction performance for six different tasks including both language generation tasks and multi-choice tasks. The x-axis represents the self-correction round and the y-axis indicates the performance evaluated on the corresponding task. The performance of self-correction improves as the interaction round progresses and converges eventually. The self-correction performance of the social bias mitigation task and the jailbreak defense task reaches the best performance in the first round and maintains this optimal performance with no modification for the rest of the interaction rounds.
The experimental results, shown in Figure 3, demonstrate the impact of self-correction across different tasks. In this figure, the x-axis represents the number of instructional rounds, while the y-axis indicates task performance. Additional experimental results are provided in Appendix A. From these results, we derive the following key observations: (1) Self-correction consistently improves performance compared to the baseline, where no self-correction instructions are employed. (2) Multi-round self-correction effectively guides LLMs towards a stable, convergent state, after which further self-correction steps do not yield significant changes in performance. (3) For multi-choice QA tasks, convergence is typically achieved after the first round, while generation tasks generally require additional rounds to reach final convergence. This disparity likely arises because free-form text generation is inherently more complex than the closed-form nature of multi-choice QA tasks.
In conclusion, the application of multi-round self-correction consistently enhances performance and eventually achieves convergence. These findings suggest that intrinsic self-correction offers convergence guarantees across a variety of tasks. In the next section, we introduce how the converged performance is related to reduced model uncertainty.
4 Model Uncertainty
In the previous section, we show empirical evidence regarding the general converged performance of intrinsic self-correction across various tasks. In this section, we provide empirical evidence showing that as model uncertainty diminishes (making LLMs less under-confident), the calibration error reduces and converges as the self-correction round progresses (for more details about model uncertainty and calibration error, please refer to Section 2). With a smaller calibration error, LLMs are more confident that their predictions are correct and aligned with the ground truth.  [23] shows that LLMs with larger model scales are well-calibrated in QA tasks since uncertainty typically reflects the model’s internal assessment on the reliability of its own responses. Building on these findings, we hypothesize that the convergence of intrinsic self-correction is driven by a reduction in uncertainty, which subsequently leads to the convergence of calibration error as the interaction rounds progress.
We adopt the method of semantic entropy [34] to estimate uncertainty for language generation tasks, which involves estimating linguistic-invariant likelihoods by the lens of semantic meanings of the text. And we utilize Rank-calibration [43] to get the calibration error for language generation tasks. Regarding multi-choice QA tasks, we consider LLMs’ predictions as a classification problem, therefore leveraging the ECE error [32], following [23]. Since the prediction logit confidence [^5] Please note higher logit confidence indicates lower uncertainty. is used as model uncertainty measurement in the ECE error, we get the normalized logits with the log-likelihoods of different choices, e.g., (a), (b), (c). We estimate model uncertainty by self-correction rounds, and pick up four social dimensions from the BBQ benchmark [8] for QA tasks.
Figure 4: The reported model uncertainty and calibration error for the language generation and QA tasks, through the lens of self-correction rounds. For QA tasks, we show results for four social bias dimensions, e.g., Physical, Sexual, Religion, and Disability. Since the ECE error converged in the first self-correction round, we add the value of baseline uncertainty and ECE error for reference, but the self-correction process starts from the first round. The uncertainty converged after 10 rounds; we show 20 rounds to indicate its convergence.
Figure 4 presents how the model uncertainty and calibration error change as the self-correction round progresses. The experimental results indicate that: (1) The uncertainty generally decreases along with more self-correction rounds across tasks. (2) All the reported tasks demonstrate a trend of converged calibration error as the rounds progress. (3) The ECE error of QA tasks converged at the first or second round, which helps to explain why the self-correction performance of QA tasks (social bias mitigation) converges in the first iteration as shown in Figure 3. (4) The RCE error of generation tasks show convergence since round 6, aligning with the trend of performance curves (text detoxification) reported in Figure 3.
The causality between model uncertainty and calibration error is bidirectional [44]. Previous studies [35, 36] demonstrate that reducing model uncertainty can help decrease calibration error by making the LLMs’ predictions more aligned with the true outcome; calibration error can also serve as a signal for the model to reassess and adjust its uncertainty. In our cases, the reduction in model uncertainty aids LLMs in achieving lower calibration error, thereby improving self-correction performance.
To summarize, during the process of intrinsic self-correction, model uncertainty consistently decreases, motivating the calibration error to diminish and eventually converge.
5 Latent Concept
In this section, we investigate how the activated latent concept evolves as the self-correction process progresses, building on the approach of identifying latent concepts to understand in-context learning [27] and the morality of LLMs [12]. In this context, a latent concept is regarded as the moral orientation underlying the input. For example, in the social bias mitigation task, the negative/immoral concept corresponds to stereotypes or discrimination, whereas the positive/moral concept represents fairness. Similarly, in the text detoxification task, concepts include toxicity and non-toxicity. We highlight two key characteristics of concepts within the context of multi-round self-correction: convergence and irreversibility. By examining these properties, we demonstrate that, when positive self-correction instructions are applied, the activated concepts consistently maintain their positive nature and eventually converge to a stable state. These characteristics offer empirical validation for the assumption underpinning the convergence of activated concepts, as discussed in Section 6.
To measure the activated concept, we employ the linear probing vector, as initially introduced by alain2016understanding [45], to interpret hidden states in black-box neural networks by training a linear classifier. The rationale behind probing vectors is to identify a space that exclusively indicates a concept, such as toxicity. For the text detoxification task, we train a toxicity classifier using a one-layer neural network on the Jigsaw dataset (further details on the probing vector can be found in Appendix B.4). We use the weight dimension of the classifier corresponding to non-toxicity as the probing vector, measuring its similarity to the hidden states across all layers and averaging the results to quantify the concept. Since social stereotypes are not explicitly stated in language but are implicitly embedded within it [46], we follow the approach of measuring concepts by constructing biased statements, as outlined by [29].
In addition to experiments demonstrating how the activated concept converges during the self-correction process in both social bias mitigation and text detoxification tasks, we conducted two additional sets of experiments to support the property of irreversibility. Specifically, we (1) introduced immoral negative instructions throughout the entire self-correction process, and (2) conducted an intervention experiment where immoral instructions were injected during rounds 2, 5, and 8 of the self-correction process. The results from these intervention experiments further underscore the strong relationship between the morality of the instructions and the moral alignment of the activated concepts. The examples of immoral instructions are shown in Appendix B.6.
Figure 5: The evolution of activated concepts for (a) QA tasks and (b) generation tasks. For the generation task, we also implement intervention experiments by injecting immoral instruction for some or all rounds.
The similarity between the activated latent concept and the probing vector across interaction rounds is presented in Figure 5. Throughout all tasks, the activation of negative concepts, such as stereotypes in QA tasks and toxicity in generation tasks, eventually converges after several rounds. Therefore, the convergence property is validated. As shown in Figure 5.(b), injecting immoral instructions results in a more toxic concept, with toxicity levels surpassing those of the baseline prompts. Conversely, when moral or immoral instructions are introduced, the resulting concept consistently converges towards being moral or immoral, respectively. Thus, the irreversibility property is validated.
We further validate the irreversibility property of activated concepts in a more challenging scenario, where the normal self-correction process is disrupted by injecting immoral instructions at specific rounds (e.g., rounds 2, 5, and 8 in our experiments shown with the red line). It is evident that once an immoral instruction is introduced, the activated concept immediately becomes significantly more toxic, even if only moral instructions were applied in previous rounds. This indicates that immoral instructions drive the activated concept towards toxicity, while moral instructions guide it towards non-toxicity. These findings strongly support the influence of the morality of the injected instructions on the morality of the activated concepts.
Our empirical analysis shows that the activated latent concept is shaped by the morality of the instruction and exhibits two key properties: convergence and irreversibility.
6 The Essential Force for Convergence